# How LLMs Learn to Reason: A Complex Network Perspective

### v2

@@ Abstract
Training large language models with Reinforcement Learning with Verifiable Rewards (RLVR) exhibits a set of distinctive and puzzling behaviors that remain poorly understood, including a two-stage learning curve, a V-shaped response-length trajectory, and a pronounced vulnerability to catastrophic forgetting. In this work, we propose that these behaviors are emergent collective phenomena governed not by neural implementation details, but by the topological evolution of the latent reasoning graph in semantic space. By demonstrating a dynamical isomorphism between a 1.5B-parameter LLM and a minimal Concept Network Model (CoNet), we trace the causal source to the self-organization of a sparse concept web pinned to an average degree of two. This geometric perspective provides a unified physical explanation for the observed anomalies: the V-shaped trajectory tracks the evolution from parallel local skill optimization to global network integration; catastrophic forgetting stems from the topological disconnection of critical “trunk” edges; and policy collapse arises from the accumulation of sequential transitions at the web’s leaf nodes, where broad exploration abruptly freezes into rigid, high-reward trajectories. Identifying a “maximally frustrated state” at the transition between learning stages, we propose Annealed-RLVR, a principled algorithm that injects a targeted SFT “heating” step to resolve this topological bottleneck. Experiments confirm that this theory-driven intervention outperforms standard RLVR on both in-distribution and out-of-distribution benchmarks (including Minerva and AIME). By recasting RLVR from black-box optimization into a predictable process of structural self-organization, our work provides a new physical intuition for engineering the emergent reasoning capabilities of future AI systems.
@@ 1 Introduction
Training large language models with reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025) has revealed four striking and still poorly understood phenomena. On math and coding benchmarks, RLVR runs typically show (i) rapid initial gains followed by a long performance plateau, even as training continues (Luo et al., 2025a; He et al., 2025; Jin et al., 2025b); (ii) a V-shaped trajectory of the length of correct solutions, where answers first become markedly shorter before later growing longer again (Luo et al., 2025a; He et al., 2025; Fatemi et al., 2025; Wang et al., 2025); (iii) pronounced vulnerability to catastrophic forgetting when RLVR-trained policies are interleaved or followed by supervised fine-tuning (SFT), with earlier reasoning skills abruptly degraded (Li & Hoiem, 2017; Luo et al., 2025b; Ding & Wang, 2025); and (iv) a collapse of solution diversity, with the policy converging onto a narrow and brittle solution manifold (He et al., 2025; Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025). Existing work offers only partial, largely phenomenological explanations for each phenomenon in isolation: plateaus and the collapse of solution diversity are linked to entropy exhaustion and over-exploitation of a fixed reward signal (Cui et al., 2025; Hao et al., 2025; Jiang et al., 2025); the initial shortening of solutions is attributed to reward-driven pruning of redundant reasoning, whereas the subsequent lengthening has been ascribed either to a shift toward solving harder instances that demand longer chains of thought, or to the emergence of multi-pass, self-reflective reasoning (“aha”-like behavior) under RLVR(Fatemi et al., 2025; Wang et al., 2025); and catastrophic forgetting is analyzed as an objective-mismatch problem between RLVR and subsequent SFT or other post-training stages (Luo et al., 2025b; Ding & Wang, 2025). However, these accounts remain fragmented: we still lack a unifying framework that ties all four behaviors to a common underlying structure or dynamical mechanism of RLVR-trained policies, and explains why they so often arise together in practice.
A growing body of work tries to explain these phenomena by modeling LLM’s reasoning process as an implicit reasoning graph—a network of semantic states and logical transitions (Wang et al., 2024; Dutta et al., 2024; Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). However, translating this idea into a fully microscopic analysis has proven extremely difficult: constructing the reasoning graph from a model’s high-dimensional latent space is a formidable open problem, preventing a direct investigation into the structural origins of the observed dynamics (See Appendix C for details).
Here, we argue that the very universality of the above RLVR dynamics provides a path forward. In complex systems, collective emergent behaviors are largely independent of microscopic details and governed by the system’s large-scale organization (Anderson, 1972; Watts & Strogatz, 1998; Barabási & Albert, 1999). This insight motivates an approach rooted in the core logic of the Renormalization Group (RG) (Wilson & Kogut, 1974): instead of directly studying the system at the full microscopic level, one can model and approach the system at a coarse-grained level to explain its macroscopic behavior. Applying this strategy, we shift our focus from the full reasoning graph at the token level to its coarse-grained backbone at the semantic level, which emerges by abstracting away fine-grained skills encoded in local network patterns like motifs or community structures (Milo et al., 2002; Fortunato, 2010; Zhang & Moore, 2014). This leads us to our central hypothesis regarding the emergent structure that forms under RLVR training for general-purpose System-2 reasoning:
Hypothesis 1.
The concept web, defined as the coarse-grained reasoning graph, is a sparse network whose effective average degree is pinned to $\left<k\right>\approx 2$.
The implications of this hypothesis are non-trivial: an average degree of $2$ implies a predominantly tree-like structure, which is efficient for generalization but inherently fragile. This topology can still support rare, large-scale cycles critical for complex operations like self-correction (as detailed in Appendix C), without significantly altering the global average degree.

### v3

# Emergent Slow Thinking in LLMs as Inverse Tree Freezing

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) enables large language models to acquire slow, multi-step reasoning from sparse final-answer signals. We provide a statistical-physics picture of this emergence. We show that an autoregressive model’s finite capacity forces it to compress its exponentially large prefix space into a Markov network of predictive states, on which slow thinking unfolds as a random walk—the Concept Network (CoNet) picture. Within CoNet, RLVR dynamics are governed by two mechanisms: merging of compatible paths and frustrated competition among incompatible ones. Together they drive the network through nucleation, growth, and freezing into multi-input, single-output directed inverse trees. The picture reproduces the training dynamics of a 1.5-billion-parameter LLM and yields three predictions: reasoning chains lengthen as a geometric necessity of sparse topology; SFT induces catastrophic forgetting through bridge-node rupture; and frustration drives policy collapse. Building on the structural timing inherent in inverse-tree freezing, we propose Annealed-RLVR—a brief SFT intervention at the moment of maximum frustration. It outperforms standard RLVR on both in- and out-of-distribution benchmarks, with the largest gains at high sampling budgets where standard RLVR collapses. The same SFT applied after the trees freeze instead triggers catastrophic forgetting, isolating timing as the active ingredient.
@@ 1 Introduction
Cognitive science draws a sharp distinction between fast thinking (System 1) and slow thinking (System 2) (Kahneman, 2011; Yang et al., 2026): the former is rapid, automatic, pattern-driven intuition; the latter is slow, deliberate, step-by-step rational reasoning. A large language model (LLM)—an autoregressive sequence generator built on the Transformer architecture (Vaswani et al., 2017)—is, by construction, a System 1 machine (Hua & Zhang, 2022): given a token sequence $(a_{1},a_{2},\ldots,a_{t})$, the model uses its parameters $\theta$ to emit, in a single forward pass, the conditional probability $\pi_{\theta}(a_{t+1}\,|\,a_{1},\ldots,a_{t})$ of the next token; the sampled $a_{t+1}$ is appended to the sequence and generation continues, with no explicit search or planning at any point. A recent finding has overturned this picture: an extraordinarily simple training paradigm—reinforcement learning with verifiable rewards (RLVR) (Liu et al., 2025; Guo et al., 2025; Team et al., 2025; Yang et al., 2025)—suffices for LLMs to spontaneously acquire slow thinking (Yang et al., 2026). After the model produces a complete response, RLVR issues a sparse $\{0,1\}$ reward based solely on the correctness of the final answer, with no supervision of intermediate steps, and the parameters $\theta$ are then updated by policy gradient. From GPT-o1 (Jaech et al., 2024) to DeepSeek-R1 (Guo et al., 2025), RLVR-trained models spontaneously unfold reasoning chains hundreds of thousands of tokens long—decomposing problems step by step, checking intermediate results, correcting errors—reaching unprecedented levels in mathematical proof and program generation (Team et al., 2026; Zeng et al., 2026; Singh et al., 2025; DeepSeek-AI, 2026). The crux is that the model is never taught how to reason—it is told only whether the answer is right—and slow thinking emerges spontaneously from the fast-thinking substrate.
From the standpoint of statistical physics, the emergence of slow thinking on top of a fast-thinking substrate carries the classical hallmark of an emergent phenomenon (Anderson, 1972; Cai et al., 2025): a separation of timescales, with microscopic degrees of freedom (individual tokens and their local statistics) unchanged yet a wholly new collective behaviour (multi-step logical reasoning) appearing at the macroscopic level. Wang et al. (Wang et al., 2025) provide direct empirical evidence for this separation through token-by-token information-entropy measurements (the model’s uncertainty over the next token) of generated sequences before and after RLVR training: about 80% of tokens lie within local computation chunks (multi-token phrases with near-deterministic transitions) where the entropy is very low (<0.5<0.5 nats) and the probability distribution barely changes between pre- and post-RLVR—these are the fast computation steps already mastered during pretraining; what RLVR genuinely reshapes is the remaining ${\sim}20\%$ of high-entropy tokens located at the transitions between chunks, which dictate the macroscopic trajectory of the reasoning chain—exactly where slow thinking takes place. Even more telling, updating only these 20% high-entropy tokens reproduces the performance gain of full training (Wang et al., 2025). Fast thinking (within chunks) stays put; slow thinking (the joints between chunks) emerges—a clean separation of timescales between the two.
Not every system trained with reinforcement learning develops emergent reasoning. AlphaGo (Silver et al., 2016), trained over the vast search space of Go ($\sim\!10^{170}$ legal positions), never spontaneously acquired a search strategy of its own—search had to be hard-coded into the algorithm via Monte Carlo tree search (MCTS) (Coulom, 2006; Browne et al., 2012). The difference lies in two structural features that LLMs possess and AlphaGo does not. The first is a domain-agnostic action-sequence representation. AlphaGo’s policy takes the current board (a state representation) as input; the move history is discarded, and a search strategy cannot be expressed in the network’s input/output format and must be bolted on externally as MCTS. An LLM’s policy, by contrast, takes the full token-generation history $(a_{1},\ldots,a_{t})$ as input (an action-sequence representation). Under this representation, fast and slow thinking share a single format—both are token sequences differing only in length—and a reasoning strategy is itself a pattern expressible and learnable within that sequence, requiring no external search module. The second is self-reference: every output $a_{t+1}$ is immediately fed back as part of the input for the next step, contributing to the prediction of $a_{t+2}$. Self-reference equips the system with an additional axis along which it can scale—inference time. Beyond the parameter count $|\theta|$ and the size of training data, it opens a third scaling channel, namely test-time scaling (Jaech et al., 2024; Snell et al., 2024; Brown et al., 2024), allowing a policy in the unified action space to deepen itself through stepwise iteration.
These two structural features explain why LLMs can develop emergent slow thinking—they possess a domain-general action space and a scalable temporal axis. They have not yet answered a more fundamental question: what does RLVR actually change at the microscopic level so that macroscopic reasoning can emerge?
The first contribution of this work is to introduce, for Transformer-class autoregressive models, an effective Markov representation built on the idea of predictive state variables (PSVs) (Littman et al., 2001). Under this representation, the exponentially large token-sequence space is compressed onto a sparse network of effective states, and reasoning is mapped to a random walk on that network. This framework also resolves a practical puzzle: why does reasoning training in LLMs require a Group Relative Policy Optimization (GRPO) based RLVR (Shao et al., 2024) while Monte Carlo tree search (MCTS) is unable to do the job? The reason is that MCTS relies on an explicit state representation (such as a board position) to construct its search tree, whereas the effective states of an action-sequence representation only emerge implicitly through training and cannot be enumerated in advance.
Building on this effective Markov representation, we propose that the training dynamics of RLVR can be mapped onto a freezing problem on a complex network. Multi-task RLVR with shared parameters merges paths and carves the high-probability edges into a sparse, multi-input, single-output directed tree—an inverse tree. The inverse tree forms spontaneously through processes analogous to nucleation and growth in a supercooled liquid (Avrami, 1939; Kelton & Greer, 2010). Almost every node on the tree has a single, near-deterministic outgoing edge, ensuring that even though reasoning may span many steps, each individual choice is nearly certain and unlikely to err—precisely the structural guarantee that allows long-range reasoning to proceed reliably. At the same time, the reasoning paths of different tasks compete on shared nodes, producing a frustration mechanism: when two reasoning trajectories share an intermediate node but demand different outgoing edges, competition at that node eventually freezes the dominant path into a deterministic transition while almost entirely erasing the subordinate one.
This theoretical framework gives a unified account of several core phenomena in RLVR training: the lengthening of reasoning chains with growing capability (Luo et al., 2025; He et al., 2025) follows naturally from the geometry of a sparse tree (Inverse Tree Picture section); the catastrophic forgetting triggered by interleaving SFT into RL (Li & Hoiem, 2017; Luo et al., 2023; Ding & Wang, 2025) stems from the rupture of critical branching nodes on the inverse tree (Catastrophic Forgetting section); and a proper understanding of the forgetting-by-frustration mechanism (Two Core Mechanisms section) points to a favourable timing window for SFT intervention—the Annealed-RLVR scheme (Annealed-RLVR section), which suppresses policy collapse (Cui et al., 2025; Yue et al., 2025) while preserving high exploration capacity.
2 Effective Markov Representation and State Compression
The autoregressive generation of a Transformer is a textbook non-Markov process: the output at step $t$ depends on the full history $(a_{1},\ldots,a_{t})$ rather than on a finite-dimensional hidden state. This dependence on the entire history is usually counted among the strengths of Transformers over models with explicit state representations such as hidden Markov models—it grants greater expressive power. By the same token, however, the dynamics of an autoregressive model is hard to analyse theoretically: treating each distinct history as a separate state yields a state space of size $V^{t}$ ($V$ the vocabulary size, $t$ the sequence length), which explodes exponentially in $t$.
The work of Littman, Sutton and Singh on predictive state representations (Littman et al., 2001) provides a key insight: for any observable dynamical system, no matter how intricate its internal mechanism, the state can be characterised equivalently by a set of predictions about future observables—the number of effective states is determined by the number of distinguishable future behaviours, not by the size of the history space. For an autoregressive model with $d=|\theta|$ parameters this means that the parameter space can encode at most $O(e^{d})$ distinguishable conditional distributions, so a great many distinct histories must necessarily map to identical future behaviour.
Following this line of thought, we define the predictive state variable (PSV): given a prompt $Q$ and an action history $h$, two pairs $(Q,h)$ and $(Q^{\prime},h^{\prime})$ belong to the same effective state if and only if they induce future conditional distributions that are indistinguishable at coarse-graining scale $\epsilon$,
$(Q,h)\sim_{\epsilon}(Q^{\prime},h^{\prime})\;\Longleftrightarrow\;\|\pi_{\theta}(\cdot\,|\,Q,h)-\pi_{\theta}(\cdot\,|\,Q^{\prime},h^{\prime})\|<\epsilon,$ (1)
where $\pi_{\theta}(\cdot\,|\,Q,h)$ denotes not the single-step distribution but the joint conditional distribution over the entire remaining token sequence. The threshold $\epsilon$ has a natural lower bound set by the finite context window and finite vocabulary. By construction, members of the same PSV class $s=[(Q,h)]_{\sim}$ produce identical future distributions up to $\epsilon$. The next token therefore depends only on $s$, not on the specific history that reached it—the defining property of a Markov state. A crucial property follows: histories from different problems can be merged into a single equivalence class through compression. This cross-problem state sharing will become the microscopic basis of path merging and inverse-tree formation in later sections.
We emphasise that the argument above establishes the existence of an effective Markov representation as a logical consequence of finite parametric capacity, but the actual partition of histories into PSV classes depends on the trained parameters and can only be constructed numerically, not derived analytically. Direct enumeration of equivalence classes in real LLMs (with $10^{9}$$10^{11}$ parameters) is currently infeasible. We therefore turn to a model system whose state space is exhaustively enumerable: the $2\times 3$ sliding puzzle ($N=120$ legal positions; the three actions $V$, $L$, $R$ are compound permutations that keep the blank in the centre column; Fig. 1a), on which we train a minimal autoregressive Transformer with GRPO (experimental details in the supplementary material). The decisive advantage of this system is that the entire branching subtree below every decision point can be expanded explicitly, so the equivalence relation (1) can be checked directly and the full time evolution of the compression ratio can be tracked.
In this experiment, a prompt $Q$ is a start state and a history $h$ is a sequence of actions $(V,L,R)$ output so far. Each prefix $(Q,h)$ constitutes a decision point at which the model must choose the next action. Most of the ${\sim}120\times 3^{12}$ possible decision points are never visited during training because the policy assigns them negligible probability; including them would inflate $R_{c}$ with spurious classes. We therefore restrict the count to $\mathcal{S}_{\text{active}}$, the set of decision points whose ancestral chain of action probabilities is $\geq 0.05$, and define the compression ratio $R_{c}=1-N_{c}/|\mathcal{S}_{\text{active}}|$, where $N_{c}$ is the number of PSV equivalence classes on $\mathcal{S}_{\text{active}}$. $R_{c}\to 1$ means that distinct histories are highly equivalent in their future behaviour and the effective state space is heavily compressed.
The model is small enough (${\sim}14$K parameters) that GRPO converges reliably from a random initialization, without pretraining. This keeps the training dynamics purely RLVR-driven, free of any confound introduced by a pretraining phase.
The temporal evolution of $R_{c}$ (Fig. 1b) traces a non-monotonic V-shape. Before training, the randomly initialized policy is near-uniform at every decision point—all nodes share essentially the same future, almost all states fall into a single PSV class, and $R_{c}$ is close to 1. Stage 1 (the rapid-learning window) is when the model begins to assign distinct action preferences to distinct states; equivalence classes split apart and $R_{c}$ plunges—this is the inevitable price of forming meaningful structure. In Stage 2, $R_{c}$ rebounds and stabilises in a high-compression regime: reasoning paths of different problems start sharing intermediate nodes, and many distinct histories collapse back onto a small number of classes—unlike the trivial initial compression, this one reflects genuine policy convergence. Throughout Stage 2, $R_{c}$ barely changes while the success rate continues to climb slowly. Visual inspection of the PSV graph (Fig. 1c) suggests that the topological skeleton is largely laid down by the end of Stage 1, with later gains arising from refinement within that already-established structure rather than from further coarse-scale reorganisation.
Figure 1: Direct measurement of PSV compression during training of a small Transformer. (a) State-graph representation of the $2\times 3$ sliding puzzle: $N=120$ legal positions form a $K=3$ regular graph; shown is a shortest solution path of distance 4. (b) Time evolution of success rate (top) and compression ratio $R_{c}$ (bottom). $R_{c}$ traces a V-shape: an initial drop driven by policy diversification, followed by a rebound and saturation in a high-compression regime (65–91%) that signals the freezing-in of the topological skeleton. (c) PSV effective state-transition graph at early (left, epoch 10) and late (right, epoch 1000) training. The number of states contracts by roughly an order of magnitude, revealing a sparse multi-input, single-output inverse-tree structure.
As a scalar, $R_{c}$ only quantifies the degree of compression and says nothing about the topology of the compressed state space. Building the effective state-transition graph—PSV equivalence classes as nodes and high-probability transitions as directed edges—makes the post-training topology directly visible (Fig. 1c): the dense, disordered initial graph contracts into a sparse multi-input, few-output directed structure—many starting paths converge into a small number of shared hub nodes and then flow, along almost unique channels, to the target.
A widely accepted view in LLM research holds that “compression is intelligence” (Delétang et al., 2023; Huang et al., 2024)—the generalization capability of a model arises fundamentally from effective compression of the training data. The state-space compression observed in our exhaustively enumerable system is therefore very likely also at work during the RLVR training of LLMs built on the same Transformer architecture. Since the state space of a real LLM cannot be directly enumerated, the next sections develop the question from both theoretical and experimental sides: how is this sparse topology sculpted by the training dynamics of RLVR, and can its formation account for the macroscopic phenomena observed in real LLM training—two-stage dynamics, growth of reasoning chains, and catastrophic forgetting?
3 The Concept Network: A Physical Picture of Slow Thinking
(a) CoNet
(b) Minimal Transformer
(c) DeepScaleR-1.5B
Figure 2: The same two-stage dynamics span three scales of model. (a) The minimal CoNet implementation: no neural network, transition probabilities as trainable parameters. (b) A 14K-parameter single-layer Transformer trained on the sliding puzzle. (c) DeepScaleR-1.5B, a 1.5-billion-parameter LLM. Top: reward; bottom: path/token length of correct responses. All three exhibit the universal Stage 1 (fast learning, shortening length) $\to$ Stage 2 (slow optimization, rebound in length) trajectory.
The PSV analysis of the previous section shows that, after RLVR training, the LLM admits an effective Markov representation. Within this representation the reasoning process is naturally a random walk on a sparse directed graph: each PSV equivalence class is a node, transition probabilities are determined implicitly by the model parameters, a single transition is one local computation (fast thinking), and a multi-step path from a question node to an answer node constitutes the full reasoning chain (slow thinking). This is the core of the Concept Network (CoNet) picture (Cai et al., 2025): slow thinking is fast thinking unfolded over time on a sparse network. This provides a different representation for RLVR from the reasoning-graph analyses discussed in (Wang et al., 2024; Minegishi et al., 2025).
In a real Transformer, every transition probability is jointly determined by self-attention, multilayer perceptrons and other intricate components; at the level of the effective Markov representation, however, all such microscopic detail is compressed into a single transition probability $p(s^{\prime}\,|\,s)$ between nodes. After pretraining, before any task-specific optimization, a great many outgoing edges at every node carry non-zero transition probability and the effective state-transition graph is a dense, high-entropy network—a random walker faces too many possible branches at every node to find a reliable path to the answer. RLVR’s role is precisely to prune these transition probabilities systematically: through GRPO and other policy-gradient methods, edges along successful paths are reinforced and the rest are suppressed; the outgoing-edge entropy of every node decreases, so that the network converges from a high-entropy uniform distribution to a low-entropy structure dominated by a few edges.
The minimal implementation of the CoNet picture treats the transition probabilities themselves as trainable parameters. We start from a $K$-regular random graph $G(N,K)$ (NN nodes, each emitting exactly $K$ directed edges); each directed edge $s\to s^{\prime}$ carries a trainable probability $p(s^{\prime}\,|\,s)$ with $\sum_{s^{\prime}}p(s^{\prime}\,|\,s)=1$; reasoning is a random walk that starts at the question node $Q$ and samples transitions according to $p(s^{\prime}\,|\,s)$, receiving reward 1 if it reaches the answer node $A$ and 0 otherwise; transition probabilities are updated by GRPO. This minimal implementation strips away every Transformer component—self-attention, MLPs, embeddings, positional encoding—and retains only the three essential ingredients of RLVR: path search on a sparse graph, shared transition probabilities, and a sparse reward at the end of the sequence.
The value of this minimal implementation is full observability: every transition probability at every node can be logged in full throughout training, and the subgraph of high-probability edges can be visualised directly. If the CoNet picture captures the core mechanism of RLVR, then the macroscopic dynamical signatures of the minimal implementation should also appear in any other system that satisfies the same conditions (sparse graph, shared parameters, sparse reward). The curve shapes depend on the initial policy and the number of training problems, and under different conditions—a pretrained model, a small problem set, or a single QA pair—they can look quite different (Power et al., 2022; Cai et al., 2025). What persists across all these regimes is cross-model consistency: the three systems track each other regardless. Figure 2 shows a representative case—a high-entropy initial policy with many training problems—for CoNet (no neural network), a 14K-parameter Transformer on the sliding puzzle, and DSR-1.5B trained under the DeepScaleR protocol (Luo et al., 2025) using the verl framework (Sheng et al., 2025; ByteDance Seed Team and verl community, ), covering five orders of magnitude in parameter count. All three display two-stage dynamics with shortening-then-lengthening responses. Among these shared features, the persistent growth of response length during Stage 2 is the most distinctive: unlike the two-stage reward profile, which is common to many optimization processes, the steady lengthening of reasoning chains is a geometric consequence of the inverse-tree picture (see below) and is not expected from generic convergence.
4 Two Core Mechanisms: Merging and Frustration
(a) Merging
(b) Frustration
(c) Step 20
(d) Step 50
(e) Step 800
[…]